# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v1

@@ 3.3 Multi-Modal Sequence Parallelism Inference Mode
The model we obtained by sequence parallelism training can handle long-context multi-modality downstream tasks. However, the most widely used inference system built on HuggingFace Transformers is usually ran on a single GPU. A lack of a distributed implementation hinders the maximum sequence length we can perform inference. The most straightforward and easy-to-use solution in HuggingFace Transformers is to use its own pipeline parallelism inference feature, which shards a single model across different devices in a layer-basis (Huang et al., 2019; Narayanan et al., 2019). However, this approach is inefficient, where only one GPU is activated at the same time. It also does not well support long sequence length because the first device needs to store more big tensors such as the inputs embedding and images, bottlenecking the whole system.
Figure 6: GPU schedule of HuggingFace pipeline versus MM-SP inference mode, demonstrated with 4 GPUs. MM-SP leverages all GPUs to jointly compute for inference results.
Thus, we implement sequence parallelism for VLMs distributed inference. Compared to the training mode, the system needs to additionally maintain tensors (e.g. input tokens and position encodings) that are progressively changing during the decoding phrase (Yu et al., 2022). In addition, the system needs to detect signals from the machine that holds the last token and accordingly terminate the distributed process. Compared to the Huggingface native pipeline parallelism strategy, our inference mode is efficient in that all devices jointly compute during the inference, accelerating the process by the number of machines times (Figure 6). In addition, it is scalable in that all memory is evenly distributed, and thus can support longer sequences with more machines.
4 Training and Data
@@ 4.1 Overall Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: context extension for LLM, Stage 4: supervised fine-tuning. Stage 1 and Stage 2 are following VILA 1.0 Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, We frozen the vision encoder and training LLM and the multi-modal projector. Afterwards, we exend the context length of LLM with text-only dataset in a continued pre-training manner. In Stage 4, we enhance the instruction following abilities by multi-task supervised fine-tuning. Both short and long data, including images, short videos and long videos related questions, are used for training. It is noted that all parameters are trainable in the final stage.
4.2 Stage1&2&3: Multimodal Alignment, Pre-training, and Short Supervised Fine-tuning
@@
Table 2: Training system throughput comparison against various Megatron-LM strategy on 32 H100 GPUs, measured in time per iteration (unit: second). ”OOM” stands for ”Out of Memory”.
@@ 4.4 Stage5: Long Supervised Fine-tuning
Long video instruction following dataset To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos with around 10 minute for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 7 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
@@ 5 Evaluation
We evaluate our full-stack solution from both system and modeling perspectives. We report our training and inference results first to verify the efficiency and scalability of our system, which makes long-context training and inference possible. We evaluate the captioning and instruction following of our long-context model.
@@ 5.1 Training and inference system
Figure 9: Maximal supported training sequence length comparison against ZigZag-RingAttn (Li et al., 2023b; Liu et al., 2024a; Zhu, 2023; Korthikanti et al., 2023) and DeepSpeed Ulysses (Jacobs et al., 2023). Our system supports higher training sequence length compared to Ulysses, and on par with ZigZag-RingAttn.
Our developed training and inference systems can be used with HuggingFace Transformer by simple monkey patching, following the popular open-source convention in (Zheng et al., 2023). In this section, we provide quantitative evaluation of training system throughput, inference system latency, and the maximal supported sequence length.
5.1.1 Training system
Baselines and hardware setup For training efficiency, we compare with a ZigZag ring-style sequence parallelism system, which is incorporates load-balanced and GPU optimization (ZigZag-RingAttn for consistency) (Li et al., 2023b; Zhu, 2023; Liu et al., 2024a; Korthikanti et al., 2023). We select a popular open-sourced implementation (Zhu, 2023). We use Fully-sharded-Data-Parallelism (FSDP) (Zhao et al., 2023) to reduce the memory occupied by models, gradients and optimizer states instead of Zero-3 (Rajbhandari et al., 2020) (Table  9). In addition, we compare with the expert-designed and highly optimized Megatron-LM (Shoeybi et al., 2019; Korthikanti et al., 2023) system. We compare with their implementation of sequence parallelism, termed “context-parallelism” (CP). We also compare with a hybrid strategy that uses tensor model parallelism (TP) within a node, and CP cross node, recommended by the Megatron-LM team as an advanced usage.
We conduct most experiments on H100 nodes, each with 8xH100 (80GB) machines equipped with intra-node NVLink, and 400 Gbps inter-node Infiniband. We conduct the maximal supported sequence length supported training experiments up to 32 A100 nodes, each equipped with 8xA100 (80GB), where the same conclusion apply for H100 because of the same total memory. We evaluate on a model of size 8B and a batch size of 1. Since Megatron-LM baseline does not natively support VLMs training and that the visual encoder is usually orders of magnitude smaller than LLMs, we provide main results for the LLM backbone, without the visual encoder. We provide an ablation of the visual encoder part in § 5.1.3.
Figure 10: Training and inference system performance obtained on 32 and 8 H100s.
[…]
Table 3: First token latency (left; unit: second) and maximal supported inference sequence length (right) comparison versus native HuggingFace pipeline.
@@ 5.1.1 Training system
Table 4: Time per iteration on the ShareGPT4Video dataset (Chen et al., 2024a) with and without our two-stage sharding strategy, measured with different settings as detailed in §$\S$ 5.1.3 (Unit: seconds).
@@ 2 GPUs 4 GPUs 8 GPUs
In summary, our training system takes the advantage of both worlds - it achieves similar scalability as ZigZag-RingAttn, and maintain the throughput of DeepSpeed-Ulysses. It achieves 1.3x speedup and a 2.5x longer context length support compared to the highly optimized Megatron-LM.
5.1.2 Inference system
We evaluate our inference system against HuggingFace Pipeline parallelism on a single node of 8 H100 GPUs and the 8B model. (Table  3). We achieve 8.2×8.2$\times$8.2 × speed up compared to HuggingFace pipeline on 8xH100 GPUs. This is mainly because Huggingface pipeline inference only activates 1 GPU at a time, where our method leverage all GPUs to jointly compute for results. Table 3 compares the maximal sequence length supported, where we find our method achieves 2.9×2.9$\times$2.9 × longer sequence length than HuggingFace pipeline. In particular, we find that during 96K inference, the largest sequence length it can support, HF pipeline stores 80GB activations on the first GPU and 18GB on later GPUs. This imbalance allocation of activation bottlenecks the maximal sequence length.
5.1.3 Effect of two-stage sharding
We ablate the effect of our two-stage sharding strategy using a video captioning dataset (Chen et al., 2024a). We compare our two-stage sharding strategy against a baseline one-stage sharding strategy that only shard by the number of images. We measure the time per iteration under different number of H100 GPUs. For $k$ GPUs, we use a total number of $k$ images per video, and a batch size of $k$. The results are shown in Table 4. We observe a speedup from 1% to 7%. The speedup comes from the long captioning, where the baseline suffers from imbalance workload without sharing by the number of text tokens.
5.2 Needle in a Haystack
5.3 General Video Understanding
@@
Table 7: Evaluation of LongVILA-Caption performance across different frame counts. LongVILA-Caption’s performance improves across all metrics. As the number of frames increases from 8 to 256, with the average score rising from 2.00 to 3.26 out of a maximum of 5. Unlike in Figure 2, the models presented in this table are both trained and evaluated using the same number of frames.
@@ 6 Conclusion
We present LongVILA, a full-stack solution for long-context visual language model, including distributed systems, model training, and data engineering. Our system can efficiently scale context length up to 2M tokens and achieve 2.1$\times$ - 5.7 $\times$ and 1.1$\times$ - 1.4$\times$ speedup than Ring Sequence Parallelism and Megatron-LM, respectively. We then meticulously curated long video datasets and train our model with a two-stage pipeline based on pre-trained VLM checkpoint. Our LongVILA-8B model extends the feasible frame number from 8 to 1024, capturing fine-grained information precisely from 2-hour video ”needle-in-a-haystack” and achieving promising results on both video QA and captioning.
@@ Appendix A Appendix
Figure 12: Convergence evaluation. We compare the training loss curve of training LongVILA with and without sequence parallelism. This figure shows convergence of LongVILA-8B model on 8 A100 GPUs with sequence parallelism degree set at 4 and pure data parallelism training. The training dataset mixture includes shot2story, ai2d and chartqa. It is obvious that the two curves align closely, indicating that our MM-SP system has no adverse impact on training quality.
@@
Table 8: Training system throughput comparison on 64 H100 GPUs, measured in time per iteration (unit: second). Ulysses does not apply here since it can only support up to 32 GPUs.
@@
Table 9: Our training system using together with FSDP (Zhao et al., 2023) or Zero-3 (Rajbhandari et al., 2020) on 32 H100 GPUs. We found that FSDP is more friendly in memory management and choose to use it in our default configuration. (Time per iteration, unit: second).

### v2

@@ 3.3 Multi-Modal Sequence Parallelism Inference Mode
The model we developed through sequence parallelism training is capable of handling long-context multi-modal downstream tasks. However, the most commonly used inference system, built on HuggingFace Transformers, typically operates on a single GPU. This lack of distributed implementation limits the maximum sequence length that can be processed during inference. The most straightforward and user-friendly solution within HuggingFace Transformers is to use its pipeline parallelism inference feature, which shards a single model across multiple devices on a layer-by-layer basis (Huang et al., 2019; Narayanan et al., 2019). However, this approach is inefficient, as it only activates one GPU at a time. Additionally, it struggles to support long sequence lengths because the first device must store large tensors, such as input embeddings and images, creating a bottleneck that hampers the entire system’s performance.
Figure 6: Inference scheduling comparison between HuggingFace Pipeline and MM-SP, illustrated with 4 GPUs. MM-SP utilizes all GPUs concurrently.
To address these limitations, we implemented sequence parallelism for distributed inference in VLMs. Unlike the training mode, the inference system must additionally manage tensors, such as input tokens and position encodings, that progressively change during the decoding phase (Yu et al., 2022). The system also needs to detect signals from the machine holding the last token to terminate the distributed process appropriately. Compared to HuggingFace’s native pipeline parallelism strategy, our inference mode is more efficient, as all devices participate in computation simultaneously, accelerating the process by a factor proportional to the number of machines (Figure 6). Furthermore, it is scalable, with memory evenly distributed across devices, enabling support for longer sequences with additional machines.
4 LongVILA Training Data and Pipeline
@@ 4.1 Overall Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: supervised fine-tuning, Stage 4: context extension for LLM, Stage 5: long supervised fine-tuning. Stage 1, 2, and 3 are following VILA Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, we frozen the vision encoder and training LLM and the multi-modal projector. In Stage 3, we fully fine-tune the model for short data instruction following, e.g., image and short video datasets. Afterwards, we extend the context length of LLM with text-only dataset in a continued pre-training manner in Stage 4. In Stage 5, we enhance the instruction following abilities by long video supervised fine-tuning. It is noted that all parameters are trainable in the final stage.
4.2 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
@@
Table 2: Training system throughput comparison across various Megatron-LM strategies on 32 H100 GPUs, measured in time per iteration (seconds). ”OOM” indicates ”Out of Memory”
@@
Figure 9: Comparison of maximum supported training sequence length with ZigZag-RingAttn (Li et al., 2023a; Liu et al., 2024a; Zhu, 2023; Korthikanti et al., 2023) and DeepSpeed-Ulysses (Jacobs et al., 2023). Our system supports longer training sequence lengths than Ulysses and is comparable to ZigZag-RingAttn.
@@ 4.4 Stage5: Long Supervised Fine-tuning
Long video instruction following dataset To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 7 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
@@ 5 Evaluation
We evaluate our full-stack solution from both system and modeling perspectives. First, we present our training and inference results to demonstrate the efficiency and scalability of our system, which enables long-context training and inference. We then assess the performance of our long-context model on captioning and instruction-following tasks.
@@ 5.1 Training and inference system
Our training and inference systems can be integrated with HuggingFace Transformers through straightforward monkey patching, in line with the popular open-source approach outlined in (Zheng et al., 2023). In this section, we present a quantitative evaluation of the training system’s throughput, the inference system’s latency, and the maximum supported sequence length.
Figure 10: Performance of training and inference systems on 32 and 8 H100 GPUs.
Table 3: Comparison of first token latency (left; seconds) and maximum supported inference sequence length (right) against the native HuggingFace Pipeline.
@@
5.1.1 Training system
Baselines and hardware setup For training efficiency, we compare our system with ZigZag ring-style sequence parallelism, which incorporates load balancing and GPU optimization (ZigZag-RingAttnfor consistency)(Li et al., 2023a; Zhu, 2023; Liu et al., 2024a; Korthikanti et al., 2023). We use a widely adopted open-source implementation (Zhu, 2023). To reduce the memory footprint of models, gradients, and optimizer states, we employ Fully-Sharded Data Parallelism (FSDP)(Zhao et al., 2023) instead of Zero-3(Rajbhandari et al., 2020) (Table 9). Additionally, we compare our system with the expert-designed and highly optimized Megatron-LM(Shoeybi et al., 2019; Korthikanti et al., 2023) system, focusing on their implementation of sequence parallelism, termed “context parallelism” (CP). We also evaluate a hybrid strategy that combines tensor model parallelism (TP) within a node and CP across nodes, as recommended by the Megatron-LM team for advanced usage.
We conduct most experiments on H100 nodes, each equipped with 8xH100 (80GB) GPUs interconnected via intra-node NVLink and 400 Gbps inter-node InfiniBand. For experiments involving the maximum supported sequence length during training, we extend the setup to 32 A100 nodes, each with 8xA100 (80GB) GPUs, where the conclusions are consistent with those for H100 due to the equivalent total memory. Our evaluations are based on an 8B model with a batch size of 1. Since the Megatron-LM baseline does not natively support VLM training and the visual encoder is typically orders of magnitude smaller than LLMs, we report the main results for the LLM backbone without the visual encoder. An ablation study of the visual encoder is provided in § 5.1.3.
Table 4: Iteration time on the ShareGPT4Video dataset (Chen et al., 2024a) with and without our two-stage sharding strategy, measured under various settings as detailed in §$\S$ 5.1.3 (seconds).
@@ 2 GPUs 4 GPUs 8 GPUs
[…]
We evaluate the maximum sequence length supported by a fixed number of GPUs by progressively increasing the per-GPU sequence length from 1k to 10k until an out-of-memory error occurs. The results are summarized in Figure 9. To ensure a fair comparison, activation checkpointing is disabled. Vanilla data parallelism fails to scale for long videos at larger cluster sizes. DeepSpeed-Ulysses partitions based on attention heads, which limits its ability to scale to higher context lengths, as the 8B model has only 32 attention heads. Consequently, our approach supports approximately 8×8$\times$8 × higher context lengths when scaled to 256 GPUs. Additionally, our system achieves a similar context length scaling as ZigZag-RingAttn, with support for over 2 million context length on 256 GPUs.
@@
In summary, our training system combines the best of both worlds—it achieves scalability comparable to ZigZag-RingAttnwhile maintaining the throughput of DeepSpeed-Ulysses. Additionally, it delivers a 1.3$\times$ speedup and supports 2.5$\times$ longer context lengths compared to the highly optimized Megatron-LM.
5.1.2 Inference system
We evaluated our inference system against HuggingFace Pipeline parallelism using a single node with 8 H100 GPUs and the 8B model (Table  3). Our system achieves an 8.2×8.2$\times$8.2 × speedup compared to HuggingFace Pipeline on 8xH100 GPUs. This significant improvement is primarily due to HuggingFace Pipeline inference activating only one GPU at a time, whereas our method leverages all GPUs to compute jointly. Table 3 compares the maximum supported sequence length, where our method supports sequences that are 2.9×2.9$\times$2.9 × longer than those supported by HuggingFace Pipeline. Specifically, during 96K sequence length inference, the maximum sequence length supported, HuggingFace Pipeline stores 80GB of activations on the first GPU and only 18GB on the remaining GPUs. This imbalanced allocation of activations limits the maximum sequence length that can be processed.
5.1.3 Effect of two-stage sharding
We evaluate the impact of our two-stage sharding strategy using a video captioning dataset (Chen et al., 2024a). We compare our two-stage sharding approach to a baseline one-stage sharding strategy that only distributes workload based on the number of images. We measure the time per iteration across different numbers of H100 GPUs. For $k$ GPUs, we use $k$ images per video and a batch size of $k$. The results, shown in Table 4, indicate a speedup ranging from 1% to 7%. This improvement is primarily observed in longer captioning tasks, where the baseline suffers from workload imbalance due to the lack of sharding based on the number of text tokens.
@@ 5.1.3 Effect of two-stage sharding
Table 7: Evaluation of LongVILA-Caption performance across different frame counts. LongVILA-Caption’s performance improves across all metrics. Unlike in Figure 2, the models presented in this table are both trained and evaluated using the same number of frames.
@@ 6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing distributed systems, model training, and data engineering. Our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism. Additionally, we have meticulously curated long video datasets and implemented a five-stage training pipeline. Our LongVILA-8B model extends the feasible frame count from 8 to 1024, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks.
@@ Appendix A Appendix
Figure 12: Convergence evaluation. We compare the training loss curves for LongVILA with and without sequence parallelism. This figure illustrates the convergence of the LongVILA-8B model on 8 A100 GPUs, with a sequence parallelism degree of 4, compared to pure data parallelism training. The training dataset includes a mixture of shot2story, ai2d, and chartqa. The two curves align closely, indicating that our MM-SP system does not negatively impact training quality.
@@
Table 8: Training system throughput comparison on 64 H100 GPUs, measured in time per iteration (seconds). Ulysses is not included in this comparison as it supports only up to 32 GPUs.
@@
Table 9: Our training system, used in conjunction with FSDP (Zhao et al., 2023) or Zero-3 (Rajbhandari et al., 2020), on 32 H100 GPUs. We found that FSDP offers more efficient memory management, which led us to select it as our default configuration. (Time per iteration, seconds).